### 15 Dec 2023 v1

Abstract
Simulating user interactions enables a more user-oriented evaluation of information retrieval (IR) systems. While user simulations are cost-efficient and reproducible, many approaches often lack fidelity regarding real user behavior. Most notably, current user models neglect the user’s context, which is the primary driver of perceived relevance and the interactions with the search results. To this end, this work introduces the simulation of context-driven query reformulations. The proposed query generation methods build upon recent Large Language Model (LLM) approaches and consider the user’s context throughout the simulation of a search session. Compared to simple context-free query generation approaches, these methods show better effectiveness and allow the simulation of more efficient IR sessions. Similarly, our evaluations consider more interaction context than current session-based measures and reveal interesting complementary insights in addition to the established evaluation protocols. We conclude with directions for future work and provide an entirely open experimental setup.
Keywords:
User Simulation Interactive Retrieval Query Generation.
1 Introduction
The Cranfield paradigm is the de facto standard approach for evaluating IR methods, allowing a fair and reproducible comparison of different retrieval systems. However, these merits come at the cost of a strong abstraction of the user behavior. The underlying user model of Cranfield-style experiments assumes that the user formulates a single query, scans the result list in its entirety until a fixed rank, and judges the relevance irrespective of earlier seen search results. In the real world, users behave differently.
Simulating user interactions offers a more cost-efficient and reproducible alternative to real-world user experiments. It allows us to conclude the generalizability of experimental results beyond the boundaries of Cranfield-style experiments. However, current endeavors lack the inclusion of the user’s context in the simulations. To this end, our work analyzes the contextual influence on simulated users, focusing on query generation.
To address the changing knowledge state of the user, we propose two new query generation methods based on generative LLMs that consider different types of context information. The methods leverage this information by incorporating it into the reformulations, thereby enhancing the fidelity of the generated queries to real-world user interactions and improving the quality of the user simulation in total.
In addition, we provide a first-of-its-kind comparison of lexical (sparse) and Transformer-based (dense) retrieval methods for simulated interactive retrieval. While dense retrieval methods typically show higher retrieval effectiveness, they come at a significantly higher cost in terms of time and resources compared to sparse methods. The simulated environment in this work allows for a direct comparison of both approaches in a user-oriented setting.
To assess the results of the simulations, we conduct an in-depth evaluation of the simulations. For a holistic understanding of the considered factors and their effects on retrieval effectiveness, the evaluation includes different perspectives of retrieval effectiveness in simulated interactive retrieval sessions. While prevalent evaluation measures report the information gain during a search session, this work also considers the effort required to acquire that knowledge beyond the reformulation of queries.
Besides the reported findings, we contribute additions for the TREC test collections in the form of new query datasets. Moreover, we provide an entirely open and reusable experimental setup. [^1] https://github.com/irgroup/SUIR
2 Related Work
Early experiments with user simulations date back to the 1980s [27, 26], but recently, the topic got more attention from the IR community [4, 5]. Like earlier works, we implement the simulations with a user model covering 1) query formulations, 2) scanning of the retrieved lists, 3) selecting and clicking appealing items, 4) reading and judging documents for relevance, and 5) inspecting other items in the result list and making stop decisions either leading to query reformulations or abandoning the search session (cf. Figure 1). The literature provides different simulation frameworks [9, 19, 23, 31]. However, this work mainly aligns with the Complex Searcher Model implemented by the SimIIR toolkit [18, 30].
Generating high-fidelity queries is an important aspect of interactive user simulations [9, 7, 10, 8]. Earlier works propose different query generation approaches that either rely on principled rules or language models. Doc2Query is a method that produces a set of questions that a given document may answer. The first presented use of these questions was to enrich documents during indexing [22]. This enrichment significantly increased the retrieval effectiveness for term-based ranking methods like BM25. [3] have shown that this method is also suitable for varying queries in a simulation environment. Recently, instruction-tuned LLMs were used to generate queries based on a given prompt providing context information about the information need [1]. This approach has not been used for simulating interactive IR sessions, which is a contribution of this work.
A commonly used measure to simulate sessions with multiple query reformulations is the Session-based DCG (sDCG[14], which adds an additional discount to the discounted cumulated gain (DCG[13] by the query position in the session. More recently, a session-based variant of rank-biased precision (RBP[20] that also builds upon a query-based discount was introduced as Session RBP (sRBP[16]. The experiments by Lipani et al. [16] suggest that sDCG and sRBP lead to different outcomes and let the authors conclude that both measures provide different evaluation perspectives. For a more in-depth review of user simulations, we refer the reader to Balog and Zhai [6].
3 Methodology
Figure 1: Focus of this work in the context of the Complex Searcher Model [19].
Similar to earlier works [9, 19, 23, 31], we ground our simulations on a defined sequence of interactions with the retrieval system and its search results, visualized in Figure 1. In this work, the particular focus is the query generation and the interaction with search engine result pages in different contexts. We evaluate simulated sessions under two different retrieval paradigms: sparse, i.e., lexical-based ranking methods, and dense, i.e., Transformer-based methods (cf. 3.1). Regarding the interaction with SERPs, we analyze user-specific click decisions and context-dependent stopping strategies (cf. 3.2). Additionally, we analyze the user’s scope of the information need and the implications of interactive feedback for query reformulations (cf. 3.3). The simulation runs we examine are analyzed with three different measures (cf. 3.4).
3.1 Retrieval Models
Table 1: Retrieval effectiveness of BM25 (sparse) and MonoT5 (dense).
Core17 Core18
Model    P@10 nDCG@10    Bpref    P@10 nDCG@10    Bpref
BM25 0.458 0.372 0.216 0.404 0.371 0.222
MonoT5@50 0.562 0.451 0.139 0.480 0.454 0.195
MonoT5@100 0.558 0.454 0.185 0.482 0.449 0.232
MonoT5@200 0.574 0.468 0.232 0.476 0.447 0.261
MonoT5@500 0.566 0.460 0.283 0.460 0.435 0.292
MonoT5@1000 0.550 0.447 0.312 0.466 0.439 0.311
The retrieval model is a key element in our simulations. In the experiments, we evaluate the benefits of using a more effective model than a baseline ranking. To this end, we compare BM25 [24] against MonoT5 [21] that is known to outperform the former method in a zero-shot setting. To better understand the quality of search results shown to the simulated users, we evaluate the retrieval effectiveness in a preliminary ad hoc retrieval experiment.
Table 1 compares BM25 to MonoT5 at different cutoffs, i.e., the cutoff level determines how many documents are retrieved by the first-stage ranker and will be re-ranked by MonoT5. In principle, the re-ranker can rely on a larger set of potentially relevant documents and place them at higher ranks. However, an increasing cutoff level does not substantially improve P@10 and nDCG@10, which can be explained by unjudged documents placed at higher ranks, as also shown by the increasing Bpref scores of MonoT5.
In our simulations, we want to compare BM25 against a strong competitor that is also computationally efficient, as the user simulations imply several other computations as well. For this reason, we consider MonoT5@100 to be sufficient for our experiments. A lower cutoff level substantially reduces the computation time and also helps to reduce the environmental impact of our research [25]. Higher cutoff levels do not lead to remarkable improvements in precision, which is of primary interest for our simulation experiments.
From a more general perspective, these evaluations further reveal the limitations of using ad-hoc test collections for simulated interactive retrieval experiments. When there are many unjudged documents in the rankings, the simulated users are indecisive about clicks, and in this case, the true potential of Transformer-based models cannot be fully analyzed in simulations when the test collection is biased towards a particular kind of retrieval method. We consider it a part of the future work to provide adequate resources, data, and tools in this regard.
3.2 Click and Stop Decisions
Table 2: Overview of our user simulation configurations based on query generation methods and click models.
(a) Query generation methods
(b) Click models
Strategy Topic Feedback Generation
GPT - - Probabilistic
GPT+ + - Probabilistic
GPT* + - Rule-based
GPT** - - Rule-based
D2Q - - Rule-based
D2Q+ - + Rule-based
D2Q++ + + Rule-based
Model rel nrel
Perfect 1.0 0.0
Navigational 0.9 0.1
Informational 0.8 0.4
Almost random 0.6 0.4
(b) Click models
To be in line with earlier work, we compare four different types of click behavior in our experiments. All of the methods are based on probabilistic decisions biased towards the level of relevance and are implemented as shown in Table 2. The perfect and almost random browsing behaviors serve as sky- and baselines, respectively. The perfect user always clicks on relevant results and never clicks non-relevant or unjudged documents, i.e., the user scents the relevant information in a strongly idealistic way. In contrast, the almost random behavior mainly neglects the relevance labels. In between, there are the informational and navigational models with different degrees of exploratory behavior. We model the stop decisions in two different ways.
Static The first stopping criterion is based on the naive assumption that the user browses a fixed number of results per page (rpp) and then continues the session by reformulating another query about the topic.
Dynamic In addition, we implement a time-based stopping criterion, first applied by Maxwell [17] in simulated IR sessions, that is based on the give-up heuristic by Krebs et al. [15] stemming from the foraging theory. More precisely, the simulated user has a fixed time budget (tnr) that starts to deplete after the last seen relevant search result. If another relevant document is seen, the time budget resets. If the budget is depleted, the user reformulates the query.
3.3 Query Generation
In the following, we describe our query generation methods, for which an overview is given in Table 2. Generally, we employ two different strategies: prompting GPT to output query strings and summarizing document contents with the help of Doc2Query (D2Q). Depending on the simulated context, we provide the simulated user with the contents of the topic file, i.e., the description and narrative, and let the user consider the feedback of earlier seen search results. On the one side, the generation methods can be classified as fully probabilistic since they rely entirely on the language models’ outputs. Conversely, there are rule-based methods that expand the query stem based on the topic’s title with newly generated query terms in a principled way.
3.3.1 Prompting query reformulations
Given the information need described by the topic file, we construct a topic-specific prompt for the LLM. More specifically, we let the LLM generate outputs based on the prompt Please generate one-hundred keyword queries about <title>. <description> <narrative>, whereas <title>, <description>, and <narrative> are taken from the respective fields of the particular topic. To better understand the effects of different prompting strategies, we consider the first generation method GPT that omits the additional topic fields <description> and <narrative> in the prompt to have context-free query variants. The second generation method GPT+ uses the entire prompt. We implement two additional methods to better understand the benefits of having different query variants with semantics. GPT* uses the topic’s title as the seed query for each query (re)formulation and expands it with a single term out of the vocabulary generated for the topic by the GPT+ strategy. Like GPT*, GPT** is also a rule-based variant based on the context-free vocabulary of GPT. For each topic in the test collections (cf. 3.5), we adapt the prompting strategy to query OpenAI’s API and parse the outputs of GPT-3.5 (more specifically, gpt-3.5-turbo-0301222https://platform.openai.com/docs/model-index-for-researchers). In total, we generated 400 queries for all 50 topics in two test collections, which resulted in a total of 40,000 generated queries that stem from the outputs of OpenAI’s LLM.
3.3.2 Doc2Query
Doc2Query relies on a set of documents to generate queries. To obtain the first set of documents, a seed query is required, which equals the <title> of the corresponding topic in this case. Our Doc2Query generation approach is based on [3] and extends the methodology by integrating background information. Unlike the GPT variants, this method allows to generate queries dynamically at runtime, taking seen results and, therefore, the context of the current session into account. To factor in the user’s changing knowledge state, terms are extracted from seen documents and added to the current knowledge state. This approach assumes that terms that come from documents actually seen are more likely to satisfy the information need than random terms from the corpus. For each query reformulation, a term is taken from the knowledge state and added to the seed query. The knowledge state at the time of the $i$-th query $KS_{i}$ is a set of terms defined according to:
$KS_{i}=\bigcup_{j\in\{1,\dots,i\}}\phi(\theta(Q_{j})).$ (1)
$\theta(Q_{j})$ is the set of seen documents for the $j$-th query. The terms extracted for a given set of documents are defined by:
$\phi(D)=\bigcup_{d\in D}\{t\in d2q(d)\,|\,idf(t)<0.5\land t\notin S\}.$ (2)
All terms identified by the Doc2Query function $d2q$ are joined, except for those with an inverse document frequency (IDF) below $0.5$ and stopwords.
To determine the effect of background information and relevance feedback on cumulative retrieval effectiveness, we present three different Doc2Query variants. D2Q serves as the baseline of the Doc2Query variants, which generates queries without background information and feedback. Here, the knowledge state is based solely on terms taken from seen documents. D2Q+ integrates the relevance feedback of the simulated user. In addition to $KS_{i}$, a knowledge state $KS_{i}^{rel}$ is used, which only contains terms from seen documents that have been marked as relevant by the simulated user. As long as $KS_{i}^{rel}$ is not empty, D2Q+ uses these terms to generate queries. Otherwise, terms from $KS_{i}$ are used:
$Q_{i+1}=\left\{\begin{array}[]{llllll}Q_{0}\cup t^{rel}&:&t^{rel}&\in&KS_{i}^{rel}&KS_{i}^{rel}\neq\emptyset\\ Q_{0}\cup t&:&t&\in&KS_{i}&else.\par\end{array}\right.$ (3)
D2Q++ extends the D2Q+ mechanism with background information. For this purpose, a set of terms is generated from the description and narrative fields. Because these terms are topic-specific, they add another form of context to the query generation. The topic-specific terms are used until no more new queries can be generated, and then the D2Q+ method takes over.
3.4 Evaluation Measures
3.4.1 Effort vs. Effect
At the most granular level, we evaluate the effect of a search session by determining the cumulated information gain (IG) over the costs of all logged interactions, i.e., the effort, by:
$\mathrm{Effect}=\sum_{s\in\mathcal{S}}\mathrm{IG}(s),\qquad\mathrm{IG}(s)=\left\{\begin{array}[]{ll}\mathrm{rel}_{d},&\mathrm{if}\ s=s_{\mathrm{rel}}\\ 0,&\mathrm{else}.\end{array}\right.$ (4)
$\mathrm{rel}_{d}$ is the relevance level of the document $d$, $\mathcal{S}$ is the set of all logged interactions, and $s_{rel}$ denotes the particular session log considering the read document as relevant. We note that in this case, the costs are modeled by several additional factors besides the query formulations, including the click decision based on the snippet, the reading of the corresponding document, and making a judgment of the document’s relevance for a given topic, and there is no discount over the progress of the session. For the costs of each action, we use the default values of SimIIR.
3.4.2 Session-based DCG
As proposed in earlier work, we evaluate the simulated sessions by the sDCG [14], which shows high correlations with real user effectiveness [11]. sDCG discounts the cumulated gain document- and also query-wise by the logarithm and the corresponding base $bq$ as well as by the query position $i$ in a session:
$\mathrm{sDCG}=\sum_{i\in\{1,\dots,n_{j}\}}\frac{\mathrm{DCG}_{q_{i}}}{1+\log_{bq}(i)},\qquad\mathrm{DCG}_{q_{i}}=\sum_{r\in\{1,\dots,n_{d}\}}\frac{2^{\mathrm{rel}_{r}}-1}{\log_{2}(r+1)},$ (5)
where $n_{j}$ denotes the number of available queries for a topic $j$ and $bq$ is a free parameter set to $4$. The DCG [13] is defined by the log-harmonic discounted sum of relevance $\mathrm{rel}_{r}$ of $n_{d}$ documents in the ranking corresponding to query $q_{i}$. Obviously, sDCG exclusively considers query formulations as costs and models the user’s stopping behavior by a log-harmonic probability distribution over the number of queries and documents.
3.4.3 Session RBP
Similar to sDCG, there is a session-based variant of RBP [20], which was introduced by Lipani et al. [16] and is defined as follows:
$\mathrm{sRBP}=(1-p)\sum_{i\in\{1,\dots,n_{j}\}}\left(\frac{p-bp}{1-bp}\right)^{i-1}\sum_{r\in\{1,\dots,n_{d}\}}(bp)^{r-1}\cdot\mathrm{rel}_{r},$ (6)
where $p$ models the user’s persistence similar to RBP and $b$ is introduced as a parameter that balances reformulating a query and continuing to browse the result list. sRBP discounts later results more than sDCG. Because of the considerable length of the simulated search sessions, we set $p$ to $0.99$ to ensure that results found later in the sessions would still contribute to the outcome. For a good trade-off between results later in the queries and results from later queries, we set $b$ to $0.9$. Like the comparison between DCG and RBP, evaluating sDCG and sRBP provides two different perspectives of retrieval effectiveness that can be explained by the underlying user models based on different probability distributions.
3.5 Implementation Details and Datasets
We implement the experimental setup with a rich set of state-of-the-art software libraries, including Pyterrier, ir_datasets, SIMIIR v2.0, and HuggingFace’s transformers. We ground the experiments on two TREC newswire test collections, i.e., the New York Times Annotated Corpus [^3] https://catalog.ldc.upenn.edu/LDC2008T19 used as part of TREC Common Core 2017 (Core17) [2], and the TREC Washington Post Corpus [^4] https://trec.nist.gov/data/wapost/ used as part of TREC Common Core 2018 (Core18) [29]. Here it is important to note that Core18 includes $\sim$26k ($\sim$4k positive) relevance labels, while Core19 includes $\sim$30k ($\sim$9k positive) labels. All of the experiments are run on a Dell workstation with an Intel i9-12900K CPU, 64GB of RAM, and an NVIDIA GeForce RTX 3070 GPU on Ubuntu 22.04 LTS.
4 Experimental Results
The following experiments investigate how the information gain develops for increasing costs. In our case, the cost is either the number of queries or the time spent, expressed in time units. The evaluations of the simulation runs are based on different user behaviors regarding click and stop decisions (cf. 3.2), two retrieval systems (cf. 3.1), and seven different methods for varying queries (cf. 3.3). Three measures (cf. 3.4) are evaluated, providing different perspectives on cost-benefit trade-offs and using two different test collections. To determine the effect of specific methods on the information gain, a default configuration is set for each method type and varied only over the aspect that is being evaluated. This default configuration is: BM25 - GPT - informational - 10rpp. In our study, we evaluate the differences between the various configurations in an exploratory manner. To make statements about the effectiveness of the different configurations, we compare both their order and the trend of the information gain pairwise over the simulated runs.
4.1 Retrieval Models and Users
This section investigates the influence of the retrieval system and different types of simulated users. Both user characteristics, i.e., the click behavior (cf. 2) and the stop decisions, are evaluated with BM25 and MonoT5. In Figure 2, we see a clear order if we look at the sDCG values for the different click profiles. The closer the user is to the perfect click behavior, the higher their information gain. Comparing all pairs of retrieval models over the click profiles (e.g., BM25-perfect vs. MonoT5-perfect) shows a slight dominance of MonoT5. This effect corresponds to the results of the ad hoc retrieval from Table 1. Compared to the sDCG plots, the effort-based evaluation shows a clear difference in user behavior, while the retrieval model plays a lesser role.
Figure 2: Results of the simulated sessions for different click behaviors. Core18 (top) and Core17 (bottom). BM25 in solid lines, MonoT5 dashed, respectively.
We distinguish two types of stop decisions (cf. 3.2). Either a new query is formulated when a fixed number of snippets has been considered (rpp) or when a defined time has passed since the last relevant document was examined (tnr). Each type is evaluated with two parameters: $tnr\in\{50,110\}$ and $rpp\in\{10,20\}$. In Figure 3, it can be seen that the dynamic click behavior positively affects the information gain. In addition, both an increase in the number of documents examined and an increase in the dynamic time budget have a positive effect. Again, the effort-based measure weights the impact of user behavior more heavily than the effectiveness of the retrieval model. Furthermore, sRBP is shown to punish dynamic stop behavior.
Figure 3: Results of the simulated sessions for different stop decisions. Core18 (top) and Core17 (bottom). BM25 in solid lines, MonoT5 dashed, respectively.
Refer to captionRefer to caption
Figure 4: Distributions for the average number of snippets examined for $i$-th query. Core18 (left) and Core17 (right).
To put the different stop profiles in an insightful context, Figure 4 shows the distribution over the average number of snippets considered across the issued queries. First, it is noticeable that both runs of rpp vary much less than the tnr variants. This effect was expected since the dynamic criterion’s stop decision depends on the relevant documents in the result list. In contrast, the upper limit is strictly fixed for the static criterion, and downward outliers only occur when the result list is smaller than the fixed value or the global time limit is reached. Furthermore, the dynamic variants examine more snippets on average at the beginning of a session and fewer snippets during the session than the corresponding static variants. For 110nr, a termination can be seen at about 80 queries. This is because the global time budget is reached here.
4.2 Query Variation
Since it has been shown in Section 4.1 that MonoT5 does not cause any substantial differences in the trends of the curves, we continue with only BM25 due to the enormous computational effort. We examine the following aspects: query generation type, topic background, and feedback (cf. Table 2). The plots of the seven different variants are shown in Figure 5.
Query generation type To determine what effect the query variation procedure has, we distinguish between probabilistic and rule-based query variations as described in Section 3.3. In our case, GPT and GPT+ stand for probabilistic variation, while the rest works ruled-based. In Figure 5, it can be clearly seen that the probabilistic variations have a strong positive effect on the information gain. All measures indicate a clear dominance of GPT and GPT+.
Topic background To examine the effect of using topic background, we compare the following strategies in pairs. GPT vs. GPT+, GPT* vs. GPT** and D2Q+ vs. D2Q++. Across all measures, the topic background has a strong positive effect on the probabilistic methods. Considering the rule-based methods (GPT* vs. GPT** and D2Q++ vs. D2Q+), the topic background does not substantially affect the information gain.
Feedback The effect of feedback on information gain is evaluated in our setting by comparing the mechanisms D2Q+ and D2Q. As described in Section 3.3, D2Q+ integrates the relevance feedback of the simulated user and accordingly favors terms that come from relevant documents. Comparing the plots displays that feedback leads to a clear increase in information gain.
Figure 5: Results of the simulated retrieval sessions for different query generation methods. Core18 (top) and Core17 (bottom).
5 Discussion
The results in Section 4.1 show that although the dense retrieval model gives slightly better results, these differences stay the same throughout the simulation. Minor differences in trends align with other studies [12, 28], and in our perspective, user behavior leads to a much greater variance in effectiveness. In particular, taking into account the context within the search session — modeled by dynamic click behavior, background knowledge, and the integration of feedback — substantially impacts information gain. From our point of view, it is crucial for future simulation experiments to integrate the user context instead of only investigating cost-intensive retrieval models.
For both test collections, the trends were similar in all evaluations, even though the absolute information gain differed. We suspect this is related to the substantial difference in the number of documents rated as relevant (cf. 3.5).
In future evaluations, the costs for specific user actions used for the effort-based measure should be specified and justified. Still, the order and trends of the effort-based evaluations resemble the sDCG evaluations closely. Furthermore, the effort-based measure can reveal effects that would remain hidden if only the number of queries were considered. For example, Figure 3 shows that the difference between 50tnr and 20rpp is more pronounced for the effort-based measure since the larger number of examined documents comes with high costs for the user that sDCG does not consider.
Regarding sRBP, sDCG also mostly supports the produced orders of the different configurations. However, the information gain stagnates more quickly, and a difference in the orders can be seen in Figure 3, so dynamic click behavior is penalized by sRBP. Both findings can be explained by the substantially stronger discounting of late results by sRBP. Interestingly, dynamic stop behavior is notably rewarded by the other measures. As Figure 4 indicates, the larger number of examined documents can likely further be attributed to this opposing evaluation behavior. Determining not only realistic durations for user actions but also estimating suitable parameters for sRBP is vital for future work to enable more accurate evaluations and could be done by analyzing real user behavior.
The use of probabilistic query generation methods shows promising results. Not only do they dominate rule-based approaches, but they are also more capable of integrating background knowledge of the topic. Integrating feedback into probabilistic approaches would be an exciting direction for future work, as our rule-based approach has clearly shown that feedback is beneficial.